OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2? y = 2 x +?

Mehek (mehek14):

plug in the ordered pair

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

?

Mehek (mehek14):

plug in 3 for x and 4 for y

Mehek (mehek14):

substitute the equation with the ordered pair

OpenStudy (anonymous):

ohh ok

Mehek (mehek14):

so your equation becomes 4 = 2(3) + b

OpenStudy (anonymous):

y=4x3?

Mehek (mehek14):

no the slope stays the same 4 is in the y's place and 3 is in the x

OpenStudy (anonymous):

oh ok

Mehek (mehek14):

like this 4 = 2(3) + b

Mehek (mehek14):

b is what you are looking for

OpenStudy (anonymous):

y=2(3)+b

OpenStudy (amorfide):

point slope form is \[y-y _{1}=m(x-x _{1})\] where m is the slope (x1,y1)=(3,4)

Mehek (mehek14):

@thebro55 no I showed you the equation looks like this: 4 = 2(3) + b

OpenStudy (anonymous):

ok

Mehek (mehek14):

now solve it

OpenStudy (anonymous):

4=6?

Mehek (mehek14):

4 = 6 + b you need to isolate b

OpenStudy (anonymous):

how?

Mehek (mehek14):

move 6 to the other side

OpenStudy (anonymous):

4=6+6

Mehek (mehek14):

no you subtract 6 on both sides

OpenStudy (anonymous):

2?

Mehek (mehek14):

4 - 6 = 6 - 6 + b

OpenStudy (anonymous):

-2=0

Mehek (mehek14):

so b = -2

OpenStudy (anonymous):

ohh. so the answer is: y=2x+-2?

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

leave out the + sign y = 2x - 2

OpenStudy (anonymous):

thx

Mehek (mehek14):

yw